# Hype, Sustainability, and the Price of the Bigger-is-Better Paradigm in AI

### v1

@@ Abstract
With the growing attention and investment in recent AI approaches such as large language models, the narrative that the larger the AI system the more valuable, powerful and interesting it is is increasingly seen as common sense. But what is this assumption based on, and how are we measuring value, power, and performance? And what are the collateral consequences of this race to ever-increasing scale? Here, we scrutinize the current scaling trends and trade-offs across multiple axes and refute two common assumptions underlying the ‘bigger-is-better’ AI paradigm: 1) that improved performance is a product of increased scale, and 2) that all interesting problems addressed by AI require large-scale models. Rather, we argue that this approach is not only fragile scientifically, but comes with undesirable consequences. First, it is not sustainable, as its compute demands increase faster than model performance, leading to unreasonable economic requirements and a disproportionate environmental footprint. Second, it implies focusing on certain problems at the expense of others, leaving aside important applications, e.g. health, education, or the climate. Finally, it exacerbates a concentration of power, which centralizes decision-making in the hands of a few actors while threatening to disempower others in the context of shaping both AI research and its applications throughout society.
@@ 1 Introduction: AI is overly focusing scale
Our field is increasingly dominated by research programs driven by a race for scale: bigger models, bigger datasets, more compute. Over the past decade, machine learning (ML) has been used to build systems used by millions, carrying out tasks such as automatic translation, news feed raking, and targeted marketing. Pursuing scale has helped improve performance of ML models on benchmarks for many tasks, with the paradigmatic example being the ability of large language models (LLMs) to encode “general knowledge” when pretrained on large amounts of text data. The perceived success of these approaches has further entrenched the assumption that bigger-is-better in AI. In this position paper, we examine the underpinnings of this assumption about scale and its collateral consequences. We argue not that scale isn’t useful, in some cases, but that there is too much focus on scale and not enough value on other research.
@@
The famous AlexNet paper (Krizhevsky et al., 2012) has been key in shaping the current era of AI, including the assumption that increased scale is the key to improved performance. While building on and acknowledging decades of work, AlexNet created the recipe for the current bigger-is-better paradigm in AI, combining GPUs, big data (at least for the time), and large-scale neural network-based approaches. The paper also demonstrated that GPUs [^1] GPUs, or Graphical Processing Units, were initially developed for video games, enabling better graphics rendering given their ability to carry out more processes in parallel compared to CPUs, or Central Processing Units, which had been the dominant hardware architecture. scale compute much better than CPUs, enabling graduate students to rig a hardware setup that produced a model which outcompeted those trained on tens of thousands of CPUs. The AlexNet authors viewed scale as a key source of their model’s benchmark beating performance: “all of our experiments suggest that our results can be improved simply by waiting for faster GPUs and bigger datasets to become available” (Krizhevsky et al., 2012, p.2). This point was further reinforced by Sutton in his “bitter lesson” article, which states that approaches based on more computation win in the long term, as the cost of of compute decreases with time (2019). The consequence of this is both an explosion in investment in large-scale AI models and a concomitant spike in the size of the notable (highly cited) models. Generative AI, whether for images or text, has taken this assumption to a new level, both within the AI research discipline and as a component of the popular ‘bigger-is-better’ narrative surrounding AI – this, of course, has corresponding training and inference compute requirements, which have ballooned in recent years (see Figure 1).
Figure 1: An explosion in model size – Top: The increase in model size means it is more and more expensive to run them in terms of RAM. Bottom: resources we need are increasing faster than available compute. Data from Epoch (2023), specific details on the plots in Appendix A.
@@
The bigger-is-better norm is also self-reinforcing, shaping the AI research field by informing what kinds of research is incentivized, which questions are asked (or remain unasked), as well as the relationship between industrial and academic actors. This is important because science, both in AI and in other disciplines, does not happen in a vacuum – it is built upon relationships and a shared pool of knowledge, in which prior work is studied, incorporated, and extended. Currently, a handful of benchmarks define how “SOTA” (state-of-the-art) is measured and understood, and in pursuit of the goal of improving performance on these benchmarks, scale has become the preferred tool for achieving progress and establishing new records. Reviewers ask for experiments at a large scale, both in the context of new models and in the context of measuring performance against existing models (e.g. fySy, 2024); scientific best practices call for running experiments many times e.g. for hyperparameter selection (Bouthillier et al., 2021). These incentives and norms contribute to pushing computing budgets beyond what is accessible to most university labs – which in turn makes many labs increasingly dependent on close ties with industry in order to secure such access (Abdalla & Abdalla, 2021; Whittaker, 2021; Abdalla et al., 2023). Taken together, we see the “bigger-is-better” norm in AI creating conditions in which it is increasingly difficult for anyone outside of large industrial labs to develop, test, and deploy SOTA AI systems.
@@
We start by discussing why this reliance on scale is misguided – we first examine this assumption in terms of how, and whether, scale leads to improvement in goal setting, finding that scale is not always correlated with better performance in certain contexts, and in fact benefits of scale tend to saturate. (section 2). Then we look at harmful consequences that result from the growth of large-scale AI. Firstly, large scale development is environmentally unsustainable (section 3), and the drive for more and more data at any cost encourages surveillance and unethical and unauditable data practices (section 4). Further, due to the expense and scarcity of core hardware and talent required to produce large-scale AI, privileging bigger-is-better increasingly concentrates power over AI in the hands of a narrow set players (section 5). We conclude by outlining how the research community can reclaim the scientific discourse in the AI field, and move away from a singular focus on scale.
@@ 2.1 Scale is one of many factors that matter
The scale of notable (highly-cited) models has massively increased over the last decade, driven by a super-exponential growth in terms of number of parameters, and amount of data and compute used (Figure 1). This growth is much more rapid than the increased capacity of hardware to execute necessary processing for model training and tuning. Indeed, while the size of large models, as measured by number of parameters, is currently doubling every 5 months (subsection A.3), the cost of 1 GB of memory has been nearly constant for a decade. This means that the resources required participate in cutting-edge AI research have increased significantly. The compute used to train an AI model went from a single day on a gaming console (in 2013) to surpassing the largest supercomputers on Earth (in 2020). A common response to concerns about the exponential growth of compute requirements is a reference to Moore’s law [^2] Moore’s law is an observation that states that the number of transistors on a microchip doubles roughly every two years. – i.e. that computational power will also increase, and this will help ensure that compute remains accessible (see Sutton, 2019). However, this is not true in practice, since compute requirements for SOTA models are surpassing improvements in computational power. For instance, Thompson et al. (2022) carried out a meta-analysis of 1,527 research papers from different ML sub-domains and extrapolated how much computational power was needed to improve upon common benchmarks like ImageNet and MS COCO. They estimate that an additional 567 times more compute would be needed to achieve a 5% error rate on ImageNet, stating that “fundamental rearchitecting is needed to lower the computational intensity so that the scaling of these problems becomes less onerous”. However, many of the architectures they survey, such as Transformer-based models, including Vision Transformers, are still the ones in use in 2024. Finally, building large AI model entails costs beyond the computational infrastructure, for instance as larger models require more human labor (Appendix C).
@@
The problem with the bigger-is-better approach is not simply that it is inaccessible. It also does not consistently produce the best models, and, after a given point, even shows diminishing returns. On many tasks, benchmark performance as a function of scale tends to saturate after a certain point (see Figure 2). In addition, there is as much variability in model performance between models within a similar size class as there is between models of different sizes (Beeching et al., 2023). Indeed, there are many factors beyond scale that are important to produce performant AI models. For instance, choosing the right model architecture for the data at hand is crucial and Transformer-based models, widely perceived to be SOTA on most or even all ML benchmarks, are not always the most fitting solution.
For instance, when working with tabular data of the kind commonly produced in enterprise environments, tree-based models produce better predictions, and and are much faster (and thus less expensive and more accessible) compared to neural network approaches. This is notably due to the fact that their inductive bias is adapted to the specificity of columnar data (Grinsztajn et al., 2022, and Figure 2a). Training or fine-tuning strategies can also play a very important role (Davidson et al., 2023) – for instance, in text embeddings where for a model of the same size, and often even for the same type of pre-trained model fine-tuned in a different manner, to produce useful similarities can markedly improve the resulting embeddings on domain-specific tasks (Muennighoff et al., 2022, and Figure 2c). We see this exemplified in the E5 model, which leverages both contrastive learning and curated data to achieve remarkable results (Wang et al., 2022). Considering text-based models, we have seen evidence that encoder-based models lead to representations that facilitate learning in relational data much better than decoder-based models, and do so in a much less compute intensive way (Grinsztajn et al., 2023).
@@ 2.2 In many applications, utility does not requires scale
g Data science                                                                                                                               Medical imaging                                                        a                                                                                                                                    b
Computer vision                                                                                c                                                                                                                                    d
Natural Language                                                                                 e                                                                                                                                    f
Figure 2: Performance as a function of scale saturates across a variety of tasks. Plots of performance as a function of scale (time or memory footprint) on benchmark data from a) tabular learning (Grinsztajn et al., 2022), b) a medical image segmentation challenge (Ma & Wang, 2023), c) computer-vision object detection (Lin et al., 2014, COCO) and d) scene parsing (Zhou et al., 2017, ADE20K), e) text embedding (Muennighoff et al., 2022), and f) text understanding (Beeching et al., 2023). Details in Appendix B.
@@
Different tasks can also call for models of different sizes, as seen on Figure 2. Using these benchmarks as a reference for comparison, we see that e.g. a 1 GB model performs well on medical image segmentation, even though the images themselves are relatively large. In computer vision, a 0.7 GB model can achieve good performance on a task such as object detection, even though scene parsing could require up to 3 GB of memory. For text processing, a natural language understanding task could be addressed with an LLM with hundreds of billions of parameters requiring more than 20 GB of memory (and multiple GPUs), even though a 1.3 GB model can also provide a good semantic embedding, which can be leveraged for solving the same task. Particularly for more focused tasks, smaller models often suffice: object detection (2c) versus scene parsing (2d), or semantics (2e) versus ’understanding’ (2f).
@@
Given this varied landscape, and the particularities in size and performance across tasks, where should we direct our research efforts? There is no single answer to this question, but it is instructive that a proposal made by Wagstaff over a decade ago called for ML with meaningful applications, each of which calls for approaches adapted to its constraints (2012), a philosophy that has become decidedly unpopular in recent years in the pursuit of ‘general-purpose’ AI models. Let us survey published applications of ML, to sheds some light on corresponding tradeoffs.
@@
If we consider health applications, where we would like machine learning to positively contribute to society, AI models are often built in data-scarce contexts on which large models can more easily overfit (Varoquaux & Cheplygina, 2022). For example, to predict epilepsy, Eriksson et al. (2023) find no prediction-performance difference between logistic regression, boosted trees, or deep learning, even as these approaches differ dramatically in terms of resources required. The largest source of medical data are probably electronic health records – Rajkomar et al. (2018) reported performance of deep learning, though their appendices show that logistic regression achieves the same performance. For education, one promise of machine learning is that it could produce customized pedagogy, shaped to the student based on which topics or skills they have learned, and which they are striving the understand. For this purpose, a random embedding of the student’s learning sequence outperforms more resource intensive optimized deep or Bayesian models (Ding & Larson, 2019). For robotics, Fu et al. (2024) present an impressively useful system that can autonomously complete many daily-life tasks such as serving dishes whose key is co-training by a human. The learning systems used are rather modest, such as a ResNet18 architecture (He et al., 2016), dating from 2016 with 11 million parameters, which only take up 22 MB of memory with a precision of bfloat16–which is dwarfed by many of today’s SOTA systems, whose size is often measured in GB (Figure 2).
@@
If we inspect applications such as those mentioned above, we can see that for applied ML research what often matters most – even if we focus on simple purpose-specific metrics – is to address well-defined and application-specific goals while respecting the compute constraints at hand. And doing so often benefits from smaller, purpose-specific approaches as opposed to larger, more generic models. The benefit of purpose-specific models is also apparent in business applications, which are, by definition, specific to given business model, and have well-defined goals (efficiency, profit, etc.) (Bernardi et al., 2019). Industry surveys indeed show that the most frequently used ML models are linear models and tree-based ones (Kumar, 2022).
@@
Happily for the AI research community, a focus on smaller models also unearths myriad unresolved scientific questions, which remain in need of scrutiny and innovative thinking independent of scale. For example, for many applications, bottlenecks lie in decision-making, calibration, and uncertainty quantification, not more data or compute as such (Van Calster et al., 2019). Similarly, causal machine learning and distributional shift are very active areas of research that are important irrespective of scale (Kaddour et al., 2022). And interpretability is still an unrealized goal for most neural-network-based approaches, even as the ability to understand and audit such systems remains centrally important in the context of meaningful real life applications, especially in high-stakes scenarios like health (Murdoch et al., 2019; Ghassemi et al., 2021).

### v2

@@ Abstract
With the growing attention and investment in recent AI approaches such as large language models, the narrative that the larger the AI system the more valuable, powerful and interesting it is is increasingly seen as common sense. But what is this assumption based on, and how are we measuring value, power, and performance? And what are the collateral consequences of this race to ever-increasing scale? Here, we scrutinize the current scaling trends and trade-offs across multiple axes and refute two common assumptions underlying the ‘bigger-is-better’ AI paradigm: 1) that performance improvements are driven by increased scale, and 2) that all interesting problems addressed by AI require large-scale models. Rather, we argue that this approach is not only fragile scientifically, but comes with undesirable consequences. First, it is not sustainable, as, despite efficiency improvements, its compute demands increase faster than model performance, leading to unreasonable economic requirements and a disproportionate environmental footprint. Second, it implies focusing on certain problems at the expense of others, leaving aside important applications, e.g. health, education, or the climate. Finally, it exacerbates a concentration of power, which centralizes decision-making in the hands of a few actors while threatening to disempower others in the context of shaping both AI research and its applications throughout society.
@@ 1 Introduction: AI is overly focusing scale
Our field is increasingly dominated by research programs driven by a race for scale: bigger models, bigger datasets, more compute. Over the past decade, machine learning (ML) has been used to build systems used by millions, carrying out tasks such as automatic translation, newsfeed raking, and targeted advertising. Pursuing scale has helped improve the performance of ML models on benchmarks for many tasks, with the paradigmatic example being the ability of large language models (LLMs) to encode “general knowledge” when pretrained on large amounts of text data. The perceived success of these approaches has further entrenched the assumption that bigger-is-better in AI. Here, we examine the underpinnings of this assumption about scale and its collateral consequences. We argue not that scale isn’t useful, in some cases, but that there is too much focus on scale and not enough value on other research.
Figure 1: An explosion in model size – Left: The increase in model size means it is more and more expensive to run them in terms of RAM. Right: resources we need are increasing faster than available compute. Data from Epoch (2023), specific details in Appendix A.
@@
The famous AlexNet paper (Krizhevsky et al., 2012) has been key in shaping the current era of AI, including the assumption that increased scale is the key to improved performance. While building on and acknowledging decades of work, AlexNet created the recipe for the current bigger-is-better paradigm in AI, combining GPUs, big data (at least for the time), and large-scale neural network-based approaches. The paper also demonstrated that GPUs [^1] GPUs, or Graphical Processing Units, were initially developed for video games, enabling better graphics rendering given their ability to carry out more processes in parallel compared to CPUs, or Central Processing Units, which had been the dominant hardware architecture. scale compute much better than CPUs, enabling graduate students to rig a hardware setup that produced a model which outcompeted those trained on tens of thousands of CPUs. The AlexNet authors viewed scale as a key source of their model’s benchmark beating performance: “all of our experiments suggest that our results can be improved simply by waiting for faster GPUs and bigger datasets to become available” (Krizhevsky et al., 2012, p.2). This point was further reinforced by Sutton in his “bitter lesson” article, which states that approaches based on more computation win in the long term, as the cost of of compute decreases with time (Sutton, 2019). With this assumption comes both an explosion in investment in large-scale AI models and a concomitant spike in the size of the notable (highly cited) models. Generative AI, whether for images or text, has taken this assumption to a new level, both within the AI research discipline and as a component of the popular ‘bigger-is-better’ narrative surrounding AI – this, of course, has impacted training and inference compute requirements, which have ballooned in recent years (see Figure 1).
@@
The bigger-is-better norm is also self-reinforcing, shaping the AI research field by informing what kinds of research is incentivized, which questions are asked (or remain unasked), as well as the relationship between industrial and academic actors. This is important because science, in AI as in other disciplines, does not happen in a vacuum – it is built upon relationships and a shared pool of knowledge, in which prior work is studied, incorporated, and extended. Currently, a handful of benchmarks define how “SOTA” (state-of-the-art) is measured and understood, and in pursuit of the goal of improving performance on these benchmarks, scale has become the preferred tool for achieving progress and establishing new records. Reviewers ask for experiments at a large scale, both in the context of new models and in the context of measuring performance against existing models (e.g. fySy, 2024); scientific best practices call for running experiments many times e.g. for hyperparameter selection (Bouthillier et al., 2021). These incentives and norms contribute to pushing computing budgets beyond what is accessible to most university labs – which in turn makes many labs increasingly dependent on close ties with industry in order to secure such access (Abdalla and Abdalla, 2021; Whittaker, 2021; Abdalla et al., 2023). Taken together, we see the “bigger-is-better” norm in AI creating conditions in which it is increasingly difficult for anyone outside of large industrial labs to develop, test, and deploy SOTA AI systems.
@@
We start by discussing why this fixation on scale is misguided – we first examine this assumption in terms of how, and whether, scale leads to improvement in goal setting, finding that scale is not well correlated with better performance in certain contexts, and in fact benefits of scale tend to saturate. (Section 2). Then we look at harmful consequences that result from the growth of large-scale AI. Firstly, large scale development is unsustainable (Section 3), and the drive for more and more data encourages unethical and unauditable data practices (Section 4). Further, due to the expense and scarcity of hardware and talent required to produce large-scale AI, bigger-is-better strategies increasingly concentrate power over AI in the hands of a narrow set players (Section 5). We conclude by outlining how the research community can reclaim the scientific discourse in the AI field, and move away from a singular focus on scale.
@@ 2.1 Scale is one of many factors that matter
Figure 2: Performance as a function of scale saturates across various tasks. Plots of performance as a function of scale (time or memory footprint) on benchmark data from a) a medical image segmentation challenge (Ma and Wang, 2023), b) computer-vision object detection (Lin et al., 2014, COCO) and c) scene parsing (Zhou et al., 2017, ADE20K), d) tabular learning (Grinsztajn et al., 2022), e) text embedding (Muennighoff et al., 2022), and f) text understanding (Beeching et al., 2023). Details in Appendix B.
@@
The scale of notable (highly-cited) models has massively increased over the last decade, driven by a super-exponential growth in terms of number of parameters, and amount of compute used (Figure 1). This growth is much more rapid than the increased capacity of hardware to execute necessary processing for model training and tuning. Indeed, while the size of large models, as measured by number of parameters, is currently doubling every 5 months (subsection A.3), the cost of 1 GB of memory has been nearly constant for a decade. This means that the resources required participate in cutting-edge AI research have increased significantly. The compute used to train an AI model went from a single day on a gaming console (in 2013) to surpassing the largest supercomputers on Earth (in 2020). A common response to concerns about the exponential growth of compute requirements is a reference to Moore’s law [^2] Moore’s law is an observation that states that the number of transistors on a microchip doubles roughly every two years. – i.e. that computational power will also increase, and this will help ensure that compute remains accessible (see Sutton, 2019). However, this is not true in practice, since compute requirements for SOTA models are surpassing improvements in computational power. For instance, Thompson et al. (2022) carried out a meta-analysis of 1,527 research papers from different ML sub-domains and extrapolated how much computational power was needed to improve upon common benchmarks like ImageNet and MS COCO. They estimated that an additional 567 times more compute would be needed to achieve a 5% error rate on ImageNet, stating that “fundamental rearchitecting is needed to lower the computational intensity so that the scaling of these problems becomes less onerous”.
Empirically, Sutton’s “bitter lesson” (Sutton, 2019) appears partly incorrect: it is not that, for AI, “general methods that leverage computation are ultimately the most effective, [because of] Moore’s law, […] continued exponentially falling cost per unit of computation”, but that increasing resources are spent on AI. This increase in resources is visible in computational costs but is also true of other costs. For instance, building larger AI models require more human labor (Appendix C).
@@
The problem with the bigger-is-better approach is not simply that it is inaccessible. It also does not consistently produce model improvements, and, after a given point, even shows diminishing returns. On many tasks, benchmark performance as a function of scale tends to saturate after a certain point (see Figure 2). In addition, there is as much variability in model performance between models within a similar size class as there is between models of different sizes (Beeching et al., 2023). Indeed, many factors beyond scale are important to produce performant AI models. Choosing the right model architecture for the data at hand is crucial. Transformer-based models, widely perceived to be SOTA on most ML benchmarks, are not always the most fitting solution. For instance, when working with tabular data of the kind commonly produced in enterprise environments, tree-based models produce better predictions, and are much faster (and thus less expensive and more accessible) compared to neural network approaches (Figure 2d). This is notably due to the fact that their inductive bias is adapted to the specificity of columnar data (Grinsztajn et al., 2022).
Even for neural networks, or attention-based models, progress has been driven by much more than scale. Training or fine-tuning strategies play a very important role (Davidson et al., 2023). For instance, in text embeddings across models of the same size, and often even for the same type of pre-trained model, fine-tuning to produce useful similarities can markedly improve the resulting embeddings on domain-specific tasks (Muennighoff et al., 2022, and Figure 2e). We see this exemplified in the E5 model, which leverages both contrastive learning and curated data to achieve remarkable results (Wang et al., 2022). Considering text-based models, encoder-based models often lead to representations that facilitate learning or language processing much better than decoder-based models, and do so in a much less compute intensive way (Grinsztajn et al., 2023; Warner et al., 2024).
@@ 2.2 In many applications, utility does not requires scale
Different tasks can also call for models of different sizes, as seen on Figure 2. Using these benchmarks as a reference for comparison, we see that e.g. a 1 GB model performs well on medical image segmentation, even though the images themselves are relatively large. In computer vision, a 0.7 GB model can achieve good performance on a task such as object detection, even though scene parsing could require up to 3 GB of memory. For text processing, a natural language understanding task could be addressed with an LLM with hundreds of billions of parameters requiring more than 20 GB of memory (and multiple GPUs), even though a 1.3 GB model can also provide a good semantic embedding, which can be leveraged for solving the same task. Particularly for more focused tasks, smaller models often suffice: object detection (2b) versus scene parsing (2c, or semantics (2e) versus ’understanding’ (2f).
@@
Given this varied landscape, and the particularities in size and performance across tasks, where should we direct our research efforts? There is no single answer to this question, but it is instructive that a proposal made by Wagstaff over a decade ago called for ML with meaningful applications, each of which calls for approaches adapted to its constraints (Wagstaff, 2012), a philosophy that has become decidedly unpopular in recent years in the pursuit of ‘general-purpose’ AI models. Let us survey published applications of ML, to sheds some light on corresponding tradeoffs. We are slowly starting to see an opening towards ML applications driven by real-world needs of end users – for instance, in the call for papers for the main track of the 2025 ICML conference, which explicitly calls for “Application-Driven Machine Learning” – a direction typically relegated to less prestigious (and less selective) workshops in other ML conferences.
@@
If we consider health applications, where we would like machine learning to positively contribute to society, AI models are often built in data-scarce contexts on which large models can more easily overfit (Varoquaux and Cheplygina, 2022). For example, to predict seizure outcome after epilepsy, Eriksson et al. (2023) find no prediction-performance difference between logistic regression, boosted trees, or deep learning, even as these approaches differ dramatically in terms of resources required. The largest source of medical data are probably electronic health records – Rajkomar et al. (2018) reported performance of deep learning, though their appendices show that logistic regression achieves the same performance. Later, Yèche et al. (2021) find deep-learning under-performing in a thorough electronic health records benchmark. For education, one promise of machine learning is that it could produce customized pedagogy, shaped to the student based on which topics or skills they have learned, and which they are striving the understand. For this purpose, a random embedding of the student’s learning sequence outperforms more resource intensive optimized deep or Bayesian models (Ding and Larson, 2019). For robotics, Fu et al. (2024) present an impressively useful system that can autonomously complete many daily-life tasks such as serving dishes whose key is co-training by a human. The learning systems used are rather modest, such as a ResNet18 architecture (He et al., 2016), dating from 2016 with 11 million parameters, which only take up 22 MB of memory with a precision of bfloat16–which is dwarfed by many of today’s SOTA systems, whose size is often measured in GB (Figure 2).
@@
If we inspect applications such as those mentioned above, we can see that for applied ML research what often matters most – even if we focus on simple purpose-specific metrics – is to address well-defined and application-specific goals while respecting the compute constraints at hand. And doing so often benefits from smaller, purpose-specific approaches as opposed to larger, more generic models. The benefit of purpose-specific models is also apparent in business applications: specificities of the business model lead to well-defined goals (efficiency, profit, etc.) (Bernardi et al., 2019). Industry surveys indeed show that the most frequently used ML models are linear models and tree-based ones (Kumar, 2022).
@@
Happily for the AI research community, a focus on smaller models also unearths myriad unresolved scientific questions, which remain in need of scrutiny and innovative thinking independent of scale. For example, for many applications, bottlenecks lie in decision-making, calibration, and uncertainty quantification (Van Calster et al., 2019), not more compute or data. Similarly, causal machine learning and distributional shift are very active areas of research that are important irrespective of scale (Kaddour et al., 2022). And interpretability is still an unrealized goal for most neural-network-based approaches, even as the ability to understand and audit such systems remains centrally important in the context of meaningful real life applications, especially in high-stakes scenarios like health (Murdoch et al., 2019; Ghassemi et al., 2021).